OpenStudy (anonymous):

what is if c is 8, then d is C-2=D

OpenStudy (jessiegonzales):

C-2=D (8)-2=D 6=D

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

8-2=6 D=6